Express the concentration of a 0.0800 M aqueous solution of fluoride, F
Answer #1

0.0800 M solution means 1.00 L of solution contains 0.0800 moles of solute.

The mass of solute is the product of the number of moles and the molar mass.

w=0.0800 \quad mol \times 18.9 \quad g/mol=1.512 \quad g

The mass of the solution is the product of its volume and density.

W= 1000 \quad mL \times 1.00 \quad g/mL=1000 \quad g

The mass percentage is the ration of the mass of solute to the mass of solution multiplied by 100.

\frac {1.512 \quadg}{1000 \quad g} \times 100=0.151 %

The concentration in ppm is the ratio of the mass of solute to the mass of solution multiplied by 1 million.

\frac {1.512 \quadg}{1000 \quad g} \times 1,000,000=1512 \quad ppm